### v1

# Longitudinal efficient adjustment sets for time-varying treatment effect estimation in nonparametric causal graphical models

Abstract
Criteria for identifying optimal adjustment sets (i.e., yielding a consistent estimator with minimal asymptotic variance) for estimating average treatment effects in parametric and nonparametric models have recently been established. In a single treatment time point setting, it has been shown that the optimal adjustment set can be identified based on a causal directed acyclic graph alone. In a longitudinal treatment setting, previous work has established graphical rules to compare the asymptotic variance of estimators based on nested time-dependent adjustment sets. However, these rules do not always permit the identification of an optimal time-dependent adjustment set based on a causal graph alone. In this paper, we extend previous results by exploiting conditional independencies that can be read from the graph. We demonstrate theoretically and empirically that our results can yield estimators with even lower asymptotic variance than those allowed by previous results. We conjecture that our new results may even allow the identification of an optimal time-dependent adjustment set based on the causal graph and provide numerical examples supporting this conjecture.
Keywords adjustment sets  $\cdot$ back-door formula  $\cdot$ causal inference  $\cdot$ nonparametric inference
1 Introduction
One of the most important challenges for causal inference when conducting observational studies lies in the ability to control for confounding. In practice, investigators often identify confounding covariates during study design and estimate the causal effect of treatment or exposure by controlling, or adjusting, for a subset of covariates using an appropriate adjustment method during data analysis. One strategy for selecting which variables to control first entails constructing a causal graph based on expert opinion and subject-matter knowledge. Then, the backdoor criterion can be used to identify sets of variables that, when adequately controlled, allow for the elimination of confounding bias [1]. Adjustment sets that allow consistent estimation of a causal effect are known as “sufficient adjustment sets”.
For a given causal graph, several sufficient adjustment sets may exist. A popular strategy for choosing among them is to select a minimal sufficient adjustment set, that is, the sufficient adjustment set that includes the fewest variables [2, 3]. Such a strategy can be appealing to proponents of parsimonious models but offers no guarantee that the resulting treatment effect estimator has a lower variance than an estimator adjusting for more variables. Alternatively, recent work has focused on the identification of the sufficient adjustment set that results in a treatment effect estimator with the lowest possible variance. Notably, Henckel et al[4] developed a criterion to identify the optimal sufficient adjustment set for estimating the effect of a point exposure on an outcome. This criterion is based on a causal graph and assumes that the relations between the variables in the graph are linear. Rotnitzky and Smucler [5] extended those results in several directions, for instance by showing that the criterion proposed by Henckel et al[4] is valid for a broad class of nonparametric estimators of the average treatment effect. They further provided a new graphical criterion for comparing adjustment sets in terms of estimation variance for estimating the effect of a time-dependent treatment subject to time-dependent confounding [5].
In this paper, we extend previous results by proposing an alternative definition of a sufficient time-dependent adjustment set and a novel criterion for comparing the variance of estimators based on different time-dependent adjustment sets. We show that employing our criterion can further decrease the variance of estimators of time-dependent treatment effects. The rest of the paper is organized as follows. In Section 2 we review some background concepts, and present the assumed data structure and the notation. In Section 3 we provide a novel definition of time-dependent adjustment sets. In Section 4 we present our new criterion and provide numerical illustrations comparing the variance of estimators based on our proposed criterion and those based on previous results. Section 5 ends with a conclusion. All proofs of results stated in the main text can be found in the Appendix.
2 Background
We begin by briefly reviewing some concepts of causal graph theory and introducing some notation. The unfamiliar reader is referred to other references for a more detailed introduction (e.g., the appendix of [6], or [1]). Let $\mathcal{G}$ be a directed acyclic graph (DAG) with nodes $\mathbf{V}$ representing the variables, that is, a graph for which all edges are unidirectional arrows and for which there is no cycle. A path is defined as a sequence of variables connected by arrows. A variable is said to be a collider in a given path if two arrows point to this variable in this path. Denote by $P$ the unknown distribution of the variables represented in the graph and by $Pa_{\mathcal{G}}(v_{k})$ the parents of $v_{k}$ according to the graph $\mathcal{G}$, that is, $Pa_{\mathcal{G}}(v_{k})$ is the set of variables in $\mathcal{G}$ that have arrows stemming from them and pointing into $v_{k}$. We say that a DAG is a causal graph if all common causes of the variables depicted in the graph are also on the graph [1] and if
$P=\prod_{v_{k}\in\mathbf{V}}f\left(v_{k}\mid Pa_{\mathcal{G}}(v_{k})\right),$
where $f(v_{k}\mid Pa_{\mathcal{G}}(v_{k}))$ is the distribution of $v_{k}$ given its parents on the graph.
Suppose that we have a study with $p+1$ follow-up times and $n$ individuals sampled from a population so that the time-ordered (or otherwise topologically ordered) data take the following form $\mathcal{O}=(\mathbf{C}_{0},A_{0},\mathbf{C}_{1},A_{1},...,\mathbf{C}_{p},A_{p},Y)$ where $\mathbf{A}=(A_{0},...,A_{p})\subset\mathbf{V}$ is a time-varying exposure or treatment, $\mathbf{C}_{k}=(C_{k,1},C_{k,2},...,C_{k,n_{ck}}),k=0,...,p$ is a vector of time-dependent covariates (with $n_{ck}$ the number of covariates at time $k$) and $Y\subset\mathbf{V}$ is the outcome of interest, measured at the end of the study. We assume that $A_{k},k=0,...,p$, correspond to finite valued random variables. For time ordered sets of variables, we use an overbar to denote the history of that variable up to a given time point. For example $\bar{\mathbf{A}}_{p}=(A_{0},...,A_{p})$. Let $Y^{\mathbf{a}}$ be the counterfactual outcome, that is, the outcome that would have been observed if, possibly contrary to fact, exposure had been $\mathbf{A}=\mathbf{a}$ where $\mathbf{a}=(a_{0},...,a_{p})$.
Based on a causal graph, it is possible to construct a single world intervention graph (SWIG) that represents the causal model (including counterfactual variables) under a hypothetical intervention [7, 8]. As will be seen shortly, we will be using these SWIGs to identify sets of variables that meet the sequential conditional exchangeability assumption of causal effects. A SWIG can be constructed from a DAG by splitting each treatment node $A_{0},...,A_{p}$ in two parts, one part for the observed and one for the value set by the (hypothetical) intervention: $A_{0}\mid a_{0},...,A_{p}\mid a_{p}$. All arrows entering into a node $A_{k}$ in the DAG also enter $A_{k}$ in the SWIG, whereas all arrows stemming from a node $A_{k}$ in the DAG now stem from $a_{k}$. Further, all descendants of a node $a_{k}$ in the SWIG are indexed by an $a_{k}$ superscript to emphasize that it is a counterfactual variable. Figure 1 presents an example of a DAG and its associated SWIG.
[…]
(a) DAG
[…]
(b) SWIG
Figure 1: A causal directed acyclic graph (DAG; a) and the corresponding single world intervention graph (SWIG; b)
Independence between variables can be read from a causal DAG or a SWIG using the concept of d-separation [1]. Consider disjoint nodes, $V_{j}$ and $V_{k}$, and set of nodes $\mathbf{V}_{\ell}$. The nodes $V_{j}$ and $V_{k}$ are d-separated by $\mathbf{V}_{\ell}$ if all paths between $V_{j}$ and $V_{k}$ are blocked given $\mathbf{V}_{\ell}$. A path is said to be blocked given $\mathbf{V}_{\ell}$ if a) it contains a non-collider that is an element of $\mathbf{V}_{\ell}$, or b) there is a collider on the path where neither it nor its descendants are in $\mathbf{V}_{\ell}$. If $V_{j}$ and $V_{k}$ are d-separated given $\mathbf{V}_{\ell}$, then $V_{j}$ and $V_{k}$ are also independent conditional on $\mathbf{V}_{\ell}$.
Our focus is on the efficient estimation of contrasts between counterfactual outcome expectations of the form $\Delta(P;\mathcal{G})\equiv\sum_{\mathbf{a}\in\mathcal{A}}c_{\mathbf{a}}\mathbb{E}_{P}(Y^{\mathbf{a}})$, where $\mathbf{a}$ is a vector taking values in the finite set $\mathcal{A}$ of all possible values that $\mathbf{A}$ can take and $c_{\mathbf{a}}\in\mathbb{R}$. For example in the context of a two time point study with a binary treatment, $\mathbf{A}=(A_{0},A_{1})$ and $\mathcal{A}=\{(0,0),(1,1),(1,0),(0,1)\}$. If $c_{(0,0)}=-1$, $c_{(1,1)}=1$, and $c_{(1,0)}=c_{(0,1)}=0$ then $\Delta(P;\mathcal{G})=\mathbb{E}_{P}(Y^{(1,1)})-\mathbb{E}_{P}(Y^{(0,0)})$. These causal contrasts can be identified from the observed data by first identifying the counterfactual expectations $\chi_{a}(P;\mathcal{G})\equiv\mathbb{E}_{P}(Y^{\mathbf{a}})$.
Definition 1.
Let ${\mathbf{Z}=(\mathbf{Z}_{0},...,\mathbf{Z}_{p})\subset\mathbf{C}}$ be time ordered, disjoint sets of variables. Similar to Rotnitzky and Smucler[5], we say that $\mathbf{Z}$ is a sufficient time-dependent adjustment set if
$\mathbb{E}_{P}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}(...\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid\bar{\mathbf{a}}_{p},\bar{\mathbf{Z}}_{p})\mid\bar{\mathbf{a}}_{p-1},\mathbf{\bar{Z}}_{p-1}\}...\mid a_{0},\mathbf{Z}_{0}])$ (1)
which can equivalently be written as
$\mathbb{E}_{P}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}\left\{\frac{I_{\bar{\mathbf{a}}}(\bar{\mathbf{A}})Y}{\prod_{k=0}^{p}\mathbb{P}(A_{k}=a_{k}\mid\bar{\mathbf{a}}_{k-1},\bar{\mathbf{Z}}_{k})}\right\},$
where $I$ is the usual indicator function.
One way to identify a sufficient time-dependent adjustment set from a causal graph is to first construct a SWIG from the causal graph, then identify a set $\mathbf{Z}_{0},...,\mathbf{Z}_{p}$ such that the sequential conditional exchangeability assumption, ${Y^{\mathbf{a}}\perp\!\!\!\perp A_{k}\mid\bar{\mathbf{A}}_{k-1},\bar{\mathbf{Z}}_{k}}$ for $k=0,...,p$, holds, where $\bar{A}_{-1}\equiv\varnothing$. Example 1 illustrates this process.
Example 1 (part 1).
Assume that Figure 1 represents a causal DAG and its associated SWIG. Using d-separation rules, it can be seen that $Y^{a}\perp\!\!\!\perp A_{0}\mid C_{0,2}$ and that $Y^{a}\perp\!\!\!\perp A_{1}^{a_{0}}\mid C_{1,2}^{a_{0}},C_{0,2},A_{0}$, for example. Under consistency (i.e., if $A_{k}=a_{k}$ then $V^{a_{k}}=V$), the second conditional independence becomes ${Y^{a}\perp\!\!\!\perp A_{1}\mid C_{1,2},C_{0,2},A_{0}}$. As such, $\mathbb{E}_{P}(Y^{\mathbf{a}})=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0},C_{0,2})\mid A_{0}=a_{0},C_{0,2}\}]$ and $\mathbf{Z}=(\mathbf{Z_{0}}=C_{0,2},\mathbf{Z_{1}}=C_{1,2})$ is a time-dependent adjustment set.
3 Proposal
We now propose an alternative definition of a sufficient time-dependent adjustment set. As Rotnitzky and Smucler[5] note, their definition of a sufficient time-dependent adjustment set ignores potential simplifications to the identification formula (1) ensuing from conditional independencies. Our alternative definition permits exploiting these conditional independencies.
Definition 2.
Let $\mathbf{Z}=(\mathbf{Z}_{0},...,\mathbf{Z}_{p})\subset\mathbf{V}\backslash Y$. We say that $\mathbf{Z}$ is a sufficient time-dependent adjustment set if
$\mathbb{E}_{P}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}(...\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid a_{p},\mathbf{Z}_{p}(\bar{\mathbf{a}}_{p-1}))\mid a_{p-1},\mathbf{Z}_{p-1}(\bar{\mathbf{a}}_{p-2})\}...\mid a_{0},\mathbf{Z}_{0}])$ (2)
or equivalently
$\mathbb{E}_{P}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}\left\{\frac{I_{\bar{\mathbf{a}}}(\bar{\mathbf{A}})Y}{\prod_{k=0}^{p}\mathbb{P}(A_{k}=a_{k}\mid\mathbf{Z}_{k}(\bar{\mathbf{a}}_{k-1}))}\right\}$
where we use the notation $\mathbf{Z}_{k}(\bar{\mathbf{a}}_{k-1})$ to emphasize that if some or all past treatments are in $\mathbf{Z}_{k}(\bar{\mathbf{a}}_{k-1})$ then they are held fixed.
Definition 2 differs from Definition 1 in a few subtle but important aspects. First, unlike Definition 1, Definition 2 allows for previous treatments to be members of $\mathbf{Z}_{k}$ (i.e., $\bar{A}_{k-1}\in\bar{\mathbf{Z}}_{k}$ is allowed), $k=1,...,p$. In addition, the expectations in the identification formula (2) are only conditional on $A_{k}$ and $\mathbf{Z}_{k}$, not on $\bar{A}_{k}$ and $\bar{\mathbf{Z}}_{k}$, $k=0,...,p$, as in Definition 1.
In practice, a sufficient time-dependent adjustment set according to Definition 2 can be constructed by first identifying a time-dependent adjustment set according to Definition 1 and exploiting conditional independencies that can be read from the DAG to exclude unnecessary variables from the identification formula. Algorithm 1 describes a process for achieving this.
Algorithm 1 Construction of a sufficient time-dependent adjustment set according to Definition 2.
1. Identify a sufficient time-dependent adjustment set $\mathbf{Z}$ according to Definition 1 using the SWIG associated with the initial DAG. Express this sufficient time-dependent adjustment set using the notation of Definition 2. In the following steps, the DAG (and not the SWIG) will be used will be used to read conditional independencies.
2. Let $Q_{p+1}=Y$.
for $k=p$ to 1 do
     3.1 Read from the DAG conditional independencies between $Q_{k+1}$ and variables in $\{A_{k},\mathbf{Z}_{k}\}$ and exclude variables from $\mathbf{Z}_{k}$ that are conditionally independent. Denote as $\mathbf{G}_{k}$ the set of variables that are retained.
     3.2 Construct a modified DAG with an additional $Q_{k}$ node and remove node $Q_{k+1}$. Add arrows pointing into $Q_{k}$ stemming from the variables $\mathbf{G}_{k}$ that were retained. Remove the node $A_{k}$ from the modified SWIG.
end for
We now return to Example 1 to illustrate this process.
Example 1 (part 2).
Step 1 of the algorithm was achieved in the first part of Example 1, where the sufficient time-dependent adjustment set $\mathbf{Z}=(\mathbf{Z_{0}}=C_{0,2},\mathbf{Z_{1}}=C_{1,2})$ was found. Using the notation introduced in Definition 2, this sufficient time-dependent adjustment set can be expressed as ${\mathbf{Z}=(\mathbf{Z}_{0}=C_{0,2},\mathbf{Z}_{1}=\{C_{1,2},A_{0},C_{0,2}\})}$, which leads to the same identification formula as previously, that is
$\mathbb{E}_{P}(Y^{\mathbf{a}})=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0},C_{0,2})\mid A_{0}=a_{0},C_{0,2}\}].$
For $k=p=1$, in Step 2, we let $Q_{2}=Y$. The resulting DAG is the same as the one in Figure 1 (a), with $Y$ replaced by $Q_{2}$. In Step 3.1, by inspecting this DAG, it can be seen that $Q_{2}\perp\!\!\!\perp C_{0,2}\mid A_{1},C_{1,2},A_{0}$ or equivalently that $Y\perp\!\!\!\perp C_{0,2}\mid A_{1},C_{1,2},A_{0}$. As such,
$\mathbb{E}_{p}(Y^{\mathbf{a}})$ $=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0},C_{0,2})\mid A_{0}=a_{0},C_{0,2}\}]$
$=\mathbb{E}_{P}[\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0})\mid A_{0}=a_{0},C_{0,2}\}]$
and we deduce that $(\mathbf{Z}_{0}=C_{0,2},\mathbf{G}_{1}=\{C_{1,2},A_{0}\})$ is a sufficient time-dependent adjustment set according to Definition 2. In Step 3.2, we produce the modified DAG in Figure 2 with $Q_{1}=\mathbb{E}_{P}(Y\mid A_{1}=a_{1},C_{1,2},A_{0}=a_{0})$ and deleted $Y$.
[…]
Figure 2: The modified causal directed acyclic graph (DAG) obtained in Step 3.2 of Example 1
For $k=0$, returning to Step 3.1, we can see from Figure 2 that $C_{0,2}$ is associated with $Q_{1}$ conditional on $A_{0}$. No additional simplification is thus possible. The resulting time-dependent adjustment set according to Definition 2 is $\mathbf{G}=(\mathbf{G}_{0}=C_{0,2},\mathbf{G}_{1}=\{C_{1,2},A_{0}\})$. Remark that this time-dependent adjustment set is not admissible according to Definition 1.
In the next section, we turn our attention to comparing the asymptotic variance of nonparametric efficient estimators of causal contrasts $\Delta(P;\mathcal{G})\equiv\sum_{\mathbf{a}\in\mathcal{A}}c_{\mathbf{a}}\mathbb{E}_{P}(Y^{\mathbf{a}})$ based on different sufficient time-dependent adjustment sets according to our proposed definition.
4 Comparison of time-dependent adjustment sets
We begin by defining some additional notation:
$b_{a_{k}}(\mathbf{Z}_{k};P)\equiv\mathbb{E}_{P}[...\mathbb{E}_{P}\{\mathbb{E}_{P}(Y\mid\mathbf{a}_{p},\mathbf{Z}_{p}(\bar{a}_{p-1}))\mid\mathbf{a}_{p-1},\mathbf{Z}_{p-1}(\bar{a}_{p-2})\}...\mid\mathbf{a}_{k},\mathbf{Z}_{k}(\bar{a}_{k-1})],$
$\pi_{a_{k}}(\mathbf{Z}_{k};P)\equiv P(A_{k}=a_{k}\mid\mathbf{Z}_{k}(\bar{a}_{k-1})),$
$\psi_{P,\mathbf{a}}(\mathbf{Z};\mathcal{G})=\dfrac{I_{\bar{\mathbf{a}}_{p}}(\bar{\mathbf{A}}_{p})}{\lambda_{\bar{\mathbf{a}}_{p}}(\bar{\mathbf{Z}}_{p};P)}\{Y-b_{\mathbf{a}_{p}}(\mathbf{Z}_{p};P)\}+\displaystyle{\sum_{k=0}^{p}}\dfrac{I_{\bar{\mathbf{a}}_{k-1}}(\bar{\mathbf{A}}_{k-1})\{b_{\mathbf{a}_{k}}(\mathbf{Z}_{k};P)-b_{\mathbf{a}_{k-1}}(\mathbf{Z}_{k-1};P)\}}{\lambda_{\bar{\mathbf{a}}_{k-1}}(\bar{\mathbf{Z}}_{k-1};P)},$
where $b_{a_{-1}}(\mathbf{Z}_{-1};P)\equiv\chi_{\mathbf{a}}(P;\mathcal{G})$, $\lambda_{\bar{\mathbf{a}}_{k-1}}(\bar{\mathbf{Z}}_{k-1};P)\equiv\displaystyle{\prod_{j=0}^{k-1}}\pi_{a_{j}}(\mathbf{Z}_{j};P)$ and $I_{\bar{\mathbf{a}}_{-1}}(\bar{\mathbf{A}}_{-1})\{\lambda_{\bar{\mathbf{a}}_{-1}}(\bar{\mathbf{Z}}_{-1};P)\}^{-1}\equiv 1$. The quantity $\psi_{P,\mathbf{a}}(\mathbf{Z};\mathcal{G})$ is the efficient influence function of the counterfactual expectation $\mathbb{E}_{P}(Y^{\mathbf{a}})$ under the nonparametric model using the time-dependent adjustment set $\mathbf{Z}$ [9]. Similar to the Cramér-Rao bound, the variance of this efficient influence function scaled by a factor $1/n$ defines a lower bound for the asymptotic variance of nonparametric estimators of $\mathbb{E}_{P}(Y^{\mathbf{a}})$ for a given time-dependent adjustment set [10, 11]. We further define $\sigma^{2}_{\mathbf{a},\mathbf{Z}}(P)$ and $\sigma^{2}_{\Delta,\mathbf{Z}}(P)$ as the asymptotic variance of a nonparametric efficient estimator of the counterfactual mean $\chi_{\mathbf{a}}(P;\mathcal{G})$ and the causal contrast $\Delta(P;\mathcal{G})$ under the time-dependent adjustment set $\mathbf{Z}$. Note that the latter can be obtained from the former using the functional delta method [12] or tricks delineated in [13]. Using this notation, the following two lemmas compare the asymptotic variance of nonparametric efficient estimators of causal contrasts between nested sufficient time-dependent adjustment sets.
Lemma 4.1 (Inclusion of additional variables $\mathbf{G}$ in a sufficient time-dependent adjustment set $\mathbf{Z}=\mathbf{B}$).
Let $\mathbf{K}=(\mathbf{K}_{0},\mathbf{K}_{1},...,\mathbf{K}_{p})$ be a sufficient time-dependent adjustment set according to Definition 1 and let ${\mathbf{B}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}}$ and ${\mathbf{G}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}\setminus\mathbf{B}_{k}},k=0,...,p$ be two disjoint subsets of $\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$ such that $(\mathbf{G}_{k},\mathbf{B}_{k})=\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$. Suppose $\mathbf{B}=(\mathbf{B}_{0},\mathbf{B}_{1},...,\mathbf{B}_{p})$ is a sufficient time-dependent adjustment set according to Definition 2 that satisfies
$A_{k}\perp\!\!\!\perp\mathbf{G}_{k}\ \mid\ \mathbf{B}_{k},\ for\ k=0,...,p.$ (3)
Then $(\mathbf{G},\mathbf{B})=\{(\mathbf{G}_{0},\mathbf{B}_{0}),(\mathbf{G}_{1},\mathbf{B}_{1}),...,(\mathbf{G}_{p},\mathbf{B}_{p})\}$ is also a sufficient time-dependent adjustment set according to Definition 2 and
$\sigma^{2}_{\mathbf{a},\mathbf{B}}(P)-\sigma^{2}_{\mathbf{a},\mathbf{G},\mathbf{B}}(P)\geq 0\ \ and\ \ \sigma^{2}_{\Delta,\mathbf{B}}(P)-\sigma^{2}_{\Delta,\mathbf{G},\mathbf{B}}(P)\geq 0.$
Lemma 4.2 (Exclusion of variables $\mathbf{B}$ from a sufficient time-dependent adjustment set $\mathbf{Z}=(\mathbf{G},\mathbf{B})$).
Let $\mathbf{K}=(\mathbf{K}_{0},\mathbf{K}_{1},...,\mathbf{K}_{p})$ be a sufficient time-dependent adjustment set according to Definition 1 and let ${\mathbf{B}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}}$ and ${\mathbf{G}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}\setminus\mathbf{B}_{k}},k=0,...,p$ be two disjoint subsets of $\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$ such that $(\mathbf{G}_{k},\mathbf{B}_{k})=\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$. Suppose $(\mathbf{G},\mathbf{B})=\{(\mathbf{G}_{0},\mathbf{B}_{0}),(\mathbf{G}_{1},\mathbf{B}_{1}),...,(\mathbf{G}_{p},\mathbf{B}_{p})\}$ is a sufficient time-dependent adjustment set according to Definition 2 that satisfies
$Y\perp\!\!\!\perp\mathbf{B}_{p}\ \mid\ \mathbf{G}_{p},\mathbf{A}_{p}$ (4)
and
$\mathbf{G}_{k}\perp\!\!\!\perp\mathbf{B}_{k-1}\ \mid\ \mathbf{G}_{k-1},\mathbf{A}_{k-1}\ for\ k=1,...,p.$ (5)
Then $\mathbf{G}=(\mathbf{G}_{0},\mathbf{G}_{1},...,\mathbf{G}_{p})$ is also a sufficient time-dependent adjustment set according to Definition 2 and
$\sigma^{2}_{\mathbf{a},\mathbf{G},\mathbf{B}}(P)-\sigma^{2}_{\mathbf{a},\mathbf{G}}(P)\geq 0\ \ and\ \ \sigma^{2}_{\Delta,\mathbf{G},\mathbf{B}}(P)-\sigma^{2}_{\Delta,\mathbf{G}}(P)\geq 0.$
The following theorem combines both of these lemmas together.
Theorem 4.3.
Let $\mathbf{K}=(\mathbf{K}_{0},\mathbf{K}_{1},...,\mathbf{K}_{p})$ be a sufficient time-dependent adjustment set according to Definition 1 and let ${\mathbf{B}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}}$ and ${\mathbf{G}_{k}\subseteq\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{j}\}\setminus\mathbf{B}_{k}},k=0,...,p$ be two disjoint subsets of $\{\bar{\mathbf{A}}_{j-1},\bar{\mathbf{K}}_{k}\}$ such that $(\mathbf{G}_{k},\mathbf{B}_{k})=\{\bar{\mathbf{A}}_{k-1},\bar{\mathbf{K}}_{k}\}$. Suppose
$\mathbf{B}=(\mathbf{B}_{0},\mathbf{B}_{1},...,\mathbf{B}_{p})$
and
$\mathbf{G}=(\mathbf{G}_{0},\mathbf{G}_{1},...,\mathbf{G}_{p})$
are two sufficient time dependent adjustment sets according to Definition 2. Suppose that
$A_{k}\perp\!\!\!\perp(\mathbf{G}_{k}\setminus\mathbf{B}_{k})\ \mid\ \mathbf{B}_{k},\ for\ k=0,...,p,$
$Y\perp\!\!\!\perp(\mathbf{B}_{p}\setminus\mathbf{G}_{p})\ \mid\ \mathbf{G}_{p},\mathbf{A}_{p}.$
and
$\mathbf{G}_{k}\perp\!\!\!\perp(\mathbf{B}_{k-1}\setminus\mathbf{G}_{k-1})\ \mid\ \mathbf{G}_{k-1},\mathbf{A}_{k-1}\ for\ k=1,...,p.$
Then for all distributions $P$ compatible with the causal graph $\mathcal{G}$
$\sigma^{2}_{\mathbf{a},\mathbf{B}}(P)-\sigma^{2}_{\mathbf{a},\mathbf{G}}(P)\geq 0\ \ and\ \ \sigma^{2}_{\Delta,\mathbf{B}}(P)-\sigma^{2}_{\Delta,\mathbf{G}}(P)\geq 0.$
We now return to Example 1 once more to illustrate how these lemmas and theorem can be used in practice to identify a time-dependent adjustment set yielding an estimator with reduced variance. Next, we revisit Example 4 of Rotnitzky and Smucler (2020) [5]. In this example, an optimal adjustment set cannot be identified based on the causal graph alone when considering Definition 1 and associated lemmas and theorem. However, our numerical example suggests that our extended results may resolve this problem.
Example 1 (part 3).
Table 1 lists all time-dependent adjustment sets according to Definition 2. Note that only adjustment sets 1-9 satisfy Definition 1. Based on Theorem 4.3, the asymptotic variance of estimators based on sets 5 and 14 can be compared, for example. Let $\mathbf{B}=(\mathbf{B_{0}}=C_{0,2},\mathbf{B_{1}}=\{C_{1,2},A_{0},C_{0,2}\})$ and $\mathbf{G}=(\mathbf{G_{0}}=C_{0,2},\mathbf{G_{1}}=\{C_{1,2},A_{0}\})$.Because $A_{1}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{1}}$, $A_{0}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{0}}$, $Y\perp\!\!\!\perp C_{0,2}\mid A_{1},C_{1,2},A_{0}$, and $(C_{1,2},C_{0})\perp\!\!\!\perp\varnothing\mid C_{0,2}A_{0}$, we know that the asymptotic variance of nonparametric efficient estimators based on set 14 is lower or equal to the asymptotic variance of nonparametric efficient estimators based on set 5. We provide a numerical illustration based on simulated data to support these theoretical results. We simulated 10,000 datasets of size $n=$ 1,000 using the following data generating equations:
$C_{01}$ $\sim TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$C_{02}$ $\sim C_{01}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$A_{0}$ $\sim Bernouilli(p=\text{expit}(C_{01}))$
$C_{11}$ $\sim A_{0}+C_{02}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$C_{12}$ $\sim C_{11}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$Y$ $\sim A_{1}+A_{0}+C_{12}+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2),$
where $TN$ refers to a truncated normal distribution (in order to avoid practical positivity violations) and $\text{expit}(\cdot)=\exp(\cdot)/(1+\exp(\cdot))$ is the inverse of the logit function. The number of replications was chosen to ensure that the Monte Carlo standard error for the standard deviation was $\approx$ 0.001, which is sufficiently small to accurately identify the adjustment set with the lowest variance. We estimated $\mathbb{E}(Y^{11})$ using a pooled longitudinal targeted maximum likelihood estimator, which is based on the nonparametric efficient influence function $\psi_{P,\mathbf{a}}(\mathbf{Z};\mathcal{G})$. This was achieved in R v4.3.2 using the ltmle package v2.0.0 and generalized linear models to model both the outcome and exposure.
Table 1: Sufficient time-dependent adjustment sets in Example 1 according to Definition 2
Adjustment set $\mathbf{Z_{0}}$ $\mathbf{Z_{1}}$ Monte Carlo
standard deviation
1 $C_{0,1}$$\{C_{1,1},A_{0},C_{0,1}\}$ 0.125
2 $C_{0,1}$$\{C_{1,2},A_{0},C_{0,1}\}$ 0.106
3 $C_{0,1}$$\{C_{1,1},C_{1,2},A_{0},C_{0,1}\}$ 0.110
4 $C_{0,2}$$\{C_{1,1},A_{0},C_{0,2}\}$ 0.118
5 $C_{0,2}$$\{C_{1,2},A_{0},C_{0,2}\}$ 0.099
6 $C_{0,2}$$\{C_{1,1},C_{1,2},A_{0},C_{0,2}\}$ 0.101
7 $\{C_{0,1},C_{0,2}\}$$\{C_{1,1},A_{0},C_{0,1},C_{0,2}\}$ 0.123
8 $\{C_{0,1},C_{0,2}\}$$\{C_{1,2},A_{0},C_{0,1},C_{0,2}\}$ 0.104
9 $\{C_{0,1},C_{0,2}\}$$\{C_{1,1},C_{1,2},A_{0},C_{0,1},C_{0,2}\}$ 0.106
10 $C_{0,1}$$\{C_{1,1},A_{0}\}$ 0.117
11 $C_{0,1}$$\{C_{1,2},A_{0}\}$ 0.099
12 $C_{0,1}$$\{C_{1,1},C_{1,2},A_{0}\}$ 0.106
13 $C_{0,2}$$\{C_{1,1},A_{0}\}$ 0.107
14 $C_{0,2}$$\{C_{1,2},A_{0}\}$ 0.088
15 $C_{0,2}$$\{C_{1,1},C_{1,2},A_{0}\}$ 0.095
16 $\{C_{0,1},C_{0,2}\}$$\{C_{1,1},A_{0},C_{0,2}\}$ 0.123
17 $\{C_{0,1},C_{0,2}\}$$\{C_{1,2},A_{0},C_{0,2}\}$ 0.104
18 $\{C_{0,1},C_{0,2}\}$$\{C_{1,1},C_{1,2},A_{0},C_{0,2}\}$ 0.106
19 $\{C_{0,1},C_{0,2}\}$$\{C_{1,1},A_{0},C_{0,1}\}$ 0.120
20 $\{C_{0,1},C_{0,2}\}$$\{C_{1,2},A_{0},C_{0,1}\}$ 0.100
21 $\{C_{0,1},C_{0,2}\}$$\{C_{1,1},C_{1,2},A_{0},C_{0,1}\}$ 0.105
22 $\{C_{0,1},C_{0,2}\}$$\{C_{1,1},A_{0}\}$ 0.112
23 $\{C_{0,1},C_{0,2}\}$$\{C_{1,2},A_{0}\}$ 0.093
24 $\{C_{0,1},C_{0,2}\}$$\{C_{1,1},C_{1,2},A_{0}\}$ 0.100
Example 2.
As a second example, we revisit Example 4 from Rotnitzky and Smucler (2020) [5]. The associated causal graph is represented in Figure 3. In this example, the optimal sufficient adjustment set cannot be uniquely identified from the causal graph alone when considering Definition 1. Indeed, as illustrated in Rotnitzky and Smucler (2020) [5], the optimal sufficient adjustment set is either 1 or 8 depending on the data-generating equations, more specifically on the relative strength of the $H\rightarrow R\rightarrow Q$ and $H\rightarrow A_{1}$ pathways.
[…]
Figure 3: Causal directed acyclic graph for Example 2 (Example 4 in Rotnitzky and Smucler, 2020)
Table 2 lists all sufficient adjustment sets according to Definition 2. Only sets 1-11 satisfy Definition 1. Using Theorem 4.3, it can be found that estimators based on set 24 have a lower variance than those based on either sets 1 and 8. For example, Theorem 4.3 allows comparing sets 1 and 24 by letting $\mathbf{B}=(\mathbf{B_{0}}=\varnothing,\mathbf{B_{1}}=\{A_{0},Q\}),\mathbf{G}=(\mathbf{G_{0}}=H,\mathbf{G_{1}}=Q)$ and noticing that $A_{1}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{1}}$, $A_{0}\perp\!\!\!\perp H\mid\varnothing$, $Y\perp\!\!\!\perp A_{0}\mid Q,A_{1}$ and $Q\perp\!\!\!\perp\varnothing\mid H,A_{0}$. Similarly, sets 8 and 24 can be compared by letting $\mathbf{B}=(\mathbf{B_{0}}=H,\mathbf{B_{1}}=\{A_{0},Q,H\}),\mathbf{G}=(\mathbf{G_{0}}=H,\mathbf{G_{1}}=Q)$ and noticing that $A_{1}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{1}}$, $A_{0}\perp\!\!\!\perp\varnothing\mid\mathbf{B_{0}}$, $Y\perp\!\!\!\perp(A_{0},H)\mid Q,A_{1}$ and $Q\perp\!\!\!\perp\varnothing\mid H,A_{0}$. To illustrate empirically these results, we generated 20,000 datasets of size $n$ = 1,000 according to the following equations
$A_{0}$ $\sim Bernoulli(p=\text{expit}(0.5))$
$H$ $\sim TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$R$ $\sim A_{0}+2.5H+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$Q$ $\sim R+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
$A_{1}$ $\sim Bernoulli(p=\text{expit}(3H))$
$Y$ $\sim A_{1}+Q+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2),$
to illustrate the case with a stronger $H\rightarrow A_{1}$ pathway, and another 20,000 datasets of size $n$ = 1,000 where the data-generating equation of $R$ was modified to
$R$ $\sim A_{0}+4H+TN(\mu=0,\sigma^{2}=1,\min=-2,\max=2)$
to illustrate the case with a stronger $H\rightarrow R\rightarrow Q$ pathway. The Monte Carlo standard error is $\leq 0.002$. It can be seen that the set with the lowest variance indeed depends on the scenario when only considering sets 1-11, but not when additionally considering sets 12-26.
Table 2: Sufficient time-dependent adjustment sets in Example 2 according to Definition 2
Adjustment set $\mathbf{Z_{0}}$ $\mathbf{Z_{1}}$ Monte Carlo SD
Stronger $H\rightarrow A_{1}$ Stronger $H\rightarrow R\rightarrow Q$
pathway pathway
1 $\varnothing$ $\{A_{0},Q\}$ 0.146 0.199
2 $\varnothing$ $\{A_{0},R\}$ 0.175 0.222
3 $\varnothing$ $\{A_{0},H\}$ 0.214 0.248
4 $\varnothing$ $\{A_{0},Q,R\}$ 0.154 0.203
5 $\varnothing$ $\{A_{0},Q,H\}$ 0.165 0.207
6 $\varnothing$ $\{A_{0},R,H\}$ 0.191 0.229
7 $\varnothing$ $\{A_{0},Q,R,H\}$ 0.165 0.207
8 $H$$\{A_{0},Q,H\}$ 0.148 0.173
9 $H$$\{A_{0},R,H\}$ 0.177 0.199
10 $H$$\{A_{0},R,Q,H\}$ 0.148 0.173
11 $H$$\{A_{0},H\}$ 0.201 0.220
12 $\varnothing$ $Q$ 0.139 0.194
13 $\varnothing$ $R$ 0.158 0.213
14 $\varnothing$ $\{Q,R\}$ 0.143 0.197
15 $\varnothing$ $\{Q,H\}$ 0.165 0.208
16 $\varnothing$ $\{R,H\}$ 0.191 0.229
17 $\varnothing$ $\{Q,R,H\}$ 0.165 0.208
18 $H$$\{Q,H\}$ 0.148 0.174
19 $H$$\{R,H\}$ 0.177 0.199
20 $H$$\{R,Q,H\}$ 0.148 0.174
21 $H$$\{A_{0},Q\}$ 0.128 0.164
22 $H$$\{A_{0},R\}$ 0.159 0.191
23 $H$$\{A_{0},R,Q\}$ 0.136 0.168
24 $H$ $Q$ 0.119 0.157
25 $H$ $R$ 0.140 0.180
26 $H$$\{R,Q\}$ 0.123 0.161
SD = Standard deviation
5 Conclusion
In this article, we built on the work of Rotnitzky and Smucler [5] and proposed an alternative definition of sufficient time-dependent adjustment sets that takes into account potential simplifications to the identification formula using conditional independencies that can be read from the causal graph. We proposed two lemmas and a theorem that allow comparing the asymptotic variance of efficient estimators based on our definition of a sufficient time-dependent adjustment sets and showed that further variance reduction can be obtained as compared to estimators based on previous results. We provided two numerical illustrations of our results. In one of those examples, the optimal adjustment set varies according to the data-generating equations for the same DAG when employing Rotnitzky and Smucler’s criterion. In contrast, among the sufficient time-dependent adjustment sets that meet our definition, the optimal set seems to remain the same regardless of the data-generating equations for the same DAG. Based on those results, we conjecture that it is possible to identify an optimal time-dependent adjustment set based on a directed acyclic graph alone when considering our definition of a sufficient time-dependent adjustment set. A formal demonstration of this result is left for future work.
Our current contribution has several important implications. For example, our results can be used directly by data-analysts when estimating the effect of a time-varying treatment in order to improve the precision of their estimates. Moreover, our results could be used in developing data-driven variable selection procedures. Indeed, several confounder selection approaches for a time-fixed covariate have been developed for targeting the optimal adjustment set (e.g.,[14, 15]). Extensions to the time-varying treatment setting may be possible based on the results of Rotnitzky and Smucler [5] and ours.
Despite the strengths of work, it is important to highlight some of its limitations. A first limitation is that we were not able to demonstrate (yet) our conjecture that our alternative definition of a sufficient time-dependent adjustment set leads to the possibility of uniquely identifying an optimal time-dependent adjustment set based on the DAG alone. A second limitation is that we have only considered DAGs where all variables are observed. In situations where some DAG vertices are unobservable, but some adjustment sets are observable, Henckel et al [4] and Rotnitzky and Smucler [5] have provided theoretical results indicating that it is not always possible to find a uniformly optimal set among observable adjustment sets.
Our current work could be extended in several directions. In studies with time-dependent confounding, it is very common for data from some individuals to be missing or incomplete due to loss to follow-up or other factors (i.e., censored data). Ignoring these losses to follow-up when identifying the adjustment set may induce selection bias. Correa et al [16] proposed a necessary and sufficient graphical criterion to estimate causal effect estimation under confounding and selection bias for the cross-sectional case. Future work could investigate whether the results of Correa et al [16] can be adapted to the longitudinal case, in particular by considering the approach presented in this paper. Finally, to facilitate the use of our proposed rules, it would be greatly valuable to develop efficient algorithms and implement them in open-source software, similar to how DAGitty can be used to identify a minimal sufficient adjustment set based on a causal graph [17], or how the optAdjSet function of the pcalg package in R can be used to identify an optimal adjustment set based on a causal graph in the single treatment time point setting.